If a faulty liquid-in-glass thermometer registers 103⁰C at steam point and has no zero error, what will it register at 60⁰C?


1
Expert's answer
2021-01-28T13:39:33-0500

Solution:


Depends on what is faulty.  If the "fault" is equal throughout (the fault is linear in behavior, which is a big assumption) then the reported value for a real value of 60 would be equal to

"\dfrac{103}{100}*60=1.03*60=61.8^o" rounded to "62^o"


I can think of two very simple potentially non-linear causes of error though.  A localized narrowing of the tube is one (non-constant change in volume with change in distance along the tube), and non-systematic unequal separation between marks (numbering ticks or lines or whatever) is another.

Answer:

"62^o"